# Multilayer Artificial Benchmark for Community Detection (mABCD)

### v1

@@ 6.2 Extracting Configuration Parameters from Real Networks
The parameter $r$ is the most challenging to retrieve, as we do not have access to the latent biscuit-like reference layer used during the network generation process. To address this, we apply a coarse estimation of this parameter. Specifically, as it was done before, we select the first layer in alphabetical order and apply the Louvain algorithm to identify its community structure, which is then treated as the reference partition. Subsequently, we detect the community structure in each layer of the network and compute the layer-specific value $r_{i}$ as the AMI between the layer’s partition and the reference partition. Since the alphabetically first layer is always used as the reference, its $r$ value will be large. Due to the non-deterministic nature of the Louvain algorithm [9], this value will not be exactly $1$, but it should be close to it.
Among the degree-oriented properties, only the retrieval of $\gamma$ is non-trivial, as it involves the general challenge of fitting an observed distribution to a specific model. To estimate this parameter, we employ the powerlaw library [1], implicitly assuming that the degree distribution follows a power-law and is not better described by any other distribution type.
The next group of parameters ($\beta$, $s$, and $S$) concerns layer-specific community structural properties. At this stage, we again employ the Louvain algorithm [9] to obtain a partition of the nodes into communities on each layer of the network. Then, a value of $\beta$ is estimated by applying the powerlaw library [1] to the sequence of community sizes. The values of $s$ and $S$ are likewise derived from this sequence.
@@
The implementation of the mABCD framework also includes parameters that do not directly reflect properties of the generated network, but instead serve as variables controlling the generation process. In the subsequent experiments, we use a fixed set of such parameters. Specifically, the dimensionality of the latent reference layer is set to $d=2$. The maximum number of iterations for sampling both degrees and community sizes is set to $1{,}000$. For the rewiring step, we use $0.05$ as the proportion of edges to be rewired in each batch, and $100$ as the total number of rewiring batches.
It is worth noting that, like any artificial model, mABCD accepts only a feasible range of configuration values. Therefore, after retrieving the parameters for timik, we tailored them to fit within the constraints of the framework. Specifically, values of $\gamma$ were limited so as not to exceed $3.0$, the minimum value of $\delta$ was set to $10$, and $s$ to $50$. Let us note that the code to extract all parameters mentioned above from any network is also present in the previously mentioned GitHub repository, so even inexperienced user can generate a “digital twin” of any real network.
@@ 6.3 Experiments
Following this hypothesis, we generated four types of networks with modified values of $\xi$ as follows: series 2 with $\xi_{i}=1.00$, series 3 with $\xi_{i}=2\xi_{i_{\textit{timik}}}$, series 4 with $\xi_{i}=0.5\xi_{i_{\textit{timik}}}$, and series 5 with $\xi_{i}=0.01$. In other words, two types of networks exhibit extreme inter-community noise levels across all layers, while the remaining two have intermediate values, determined proportionally to those of the baseline. We then selected a spreading regime for the baseline networks (i.e., “digital twins” of timik, denoted as series 1). Specifically, we used the Neighbourhood Size Discount [18] to select seeds, and considered two sets of parameters for each protocol function: $(\dot{\delta},\dot{\pi},s)\in\{AND\}\times\{0.15,0.2,0.25,0.3\}\times\{1,3,5,7\}$ and $(\dot{\delta},\dot{\pi},s)\in\{OR\}\times\{0.03,0.05,0.07,0.09\}\times\{1,5,10,15\}$, where $\dot{\delta}$ and $\dot{\pi}$ denote the parameters of the MICM (see Section 6.1), while $s$ refers to the seed set budget, expressed as a percentage of the number of actors constituting a given network.
@@
With the second experiment, we aimed to address one of the fundamental questions in the domain of social network analysis: is it possible to compress a network in such a way that it preserves its topology while being smaller? Our goal was to examine this issue in the context of spreading phenomena. Specifically, we were interested in whether manipulating the number of actors, while preserving all other structural properties of the network (as dictated by the configuration model underlying mABCD), would result in different diffusion characteristics. It is worth noting that the approach adopted in this study can be considered novel, as most existing work reduces graphs using dedicated compression techniques rather than by explicitly employing standalone graph generators (for a comprehensive review, see [52]).
@@
As one can note, for both protocol functions and across all evaluated activation probabilities, downsampling the network does not disturb spreading effectiveness (the differences are minimal). From the other hand, when the network gets upsampled, we observe slight, but non-negligible changes, in the diffusion — the more upsampled the network is, the higher $\Gamma$ change we observe. However, the disparities are not as big as in the results from Section 6.3. At this point, one might wonder why network downsampling does not alter the spreading trend, while upsampling does. A possible explanation lies in the average degree of the evaluated networks, as this parameter has a substantial impact on the effectiveness of diffusion under the MICM. We observe that downsampling has little effect on the average degree, whereas upsampling increases it by approximately $0.9$ (from around $20.2$ to $21.1$). Consequently, in the upsampled graphs, the diffusion process may be more effective due to this slight increase in connectivity, especially when the seed set is selected using a heuristic that prioritizes central actors rather than selects them at random.
@@
As in the previous experiments, the networks were also generated according to the configuration derived from timik. All of them had the $r$ value on the first layer set to $1.000$ in order to comply with the assumed seed selection scenario. However, they differed in the $r$ values on the remaining layers, as follows. Two series were generated to represent stronger community correlations: series 10 with $r_{2}=r_{3}=1.000$, and series 11 with $r_{2}=r_{3}=0.667$. Additionally, networks with weaker correlations were generated: series 12 with $r_{2}=r_{3}=0.333$, and series 13 with $r_{2}=r_{3}=0.001$. The spreading conditions applied were the same as in the previous experiments, with one difference in the seed selection method. Specifically, instead of employing a heuristic that considers actors’ representations across all network layers, we modified it to consider only the first layer (which is maximally correlated with the latent layer used in network generation). Consequently, the Neighbour Size Discount heuristic was reduced to its baseline version designed for single-layer networks: Degree Centrality Discount [16] and with that method a seed set was selected.
@@ 7 Conclusions
In this work, we have presented mABCD, a framework for generating multilayer networks with community structure and inter-layer relationships. We hope it addresses the existing gap in the field through its highly customizable configuration options and efficient algorithmic design. Furthermore, we release its implementation in Julia, along with Python ports, to make mABCD accessible to the broader research community. As mABCD builds upon the well-established generator introduced in [34], it can be regarded as a new member of the ABCD family. In addition to outlining the essential components of the model, we have validated its properties, demonstrating that mABCD produces stable results and outperforms its closest competitors in both efficiency and usage flexibility.
The paper also shows potential applications of the highly customizable network generation process, illustrated in the context of studying spreading phenomena. Through three experiments, we examined whether and how selected parameters of the mABCD configuration model contribute to diffusion effectiveness. The results are nuanced rather than clear-cut. While there is a consistent trend with respect to the number of actors, the level of inter-community noise affects diffusion depending on the imposed regime. Finally, the strength of correlation between groups appears to have little influence on information spread. Altogether, these findings suggest that the proposed framework is capable of supporting the analysis of spreading processes in multilayer networks, albeit up to certain extend.
Future work may proceed in two directions. The first concerns the ABCD model itself, which could be extended to other types of networks, such as temporal ones. The second relates to the experiments involving spreading phenomena. In this context, it would be valuable to evaluate how other parameters of mABCD influence the dynamics of spreading processes. However, a more important and promising direction is the development of a metric for quantifying the difference between a real-world network and its “digital twin” generated by mABCD. Such a method could open a new avenue for influence maximization studies, as it would allow results and conclusions to be more directly related to real-world systems.
@@
This research was partially supported by the National Science Centre, Poland, under Grant no. 2022/45/B/ST6/04145888https://multispread.pwr.edu.pl/, the Polish Ministry of Education and Science within the programme “International Projects Co-Funded”, and the EU under the Horizon Europe, grant no. 101086321 OMINO. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the National Science Centre, Polish Ministry of Education and Science, the EU or the European Research Executive Agency.

### v2

@@ 6.2 Extracting configuration parameters from real networks
The parameter $r$ is the most challenging to retrieve, as we do not have access to the latent biscuit-like reference layer used during the network generation process. To address this, we apply a coarse estimation of this parameter. Specifically, as it was done before, we select the first layer in alphabetical order and apply the Louvain algorithm to identify its community structure, which is then treated as the reference partition. Subsequently, we detect the community structure in each layer of the network and compute the layer-specific value $r_{i}$ as the AMI between the layer’s partition and the reference partition. Since the alphabetically first layer is always used as the reference, its $r$ value will be large. Due to the non-deterministic nature of the Louvain algorithm [blondel2008fast], this value will not be exactly $1$, but it should be close to it.
Among the degree-oriented properties, only the retrieval of $\gamma$ is non-trivial, as it involves the general challenge of fitting an observed distribution to a specific model. To estimate this parameter, we employ the powerlaw library [alstott2014powerlaw], implicitly assuming that the degree distribution follows a power-law and is not better described by any other distribution type.
The next group of parameters ($\beta$, $s$, and $S$) concerns layer-specific community structural properties. At this stage, we again employ the Louvain algorithm [blondel2008fast] to obtain a partition of the nodes into communities on each layer of the network. Then, a value of $\beta$ is estimated by applying the powerlaw library [alstott2014powerlaw] to the sequence of community sizes. The values of $s$ and $S$ are likewise derived from this sequence.
@@
The implementation of the mABCD framework also includes parameters that do not directly reflect properties of the generated network, but instead serve as variables controlling the generation process. In the subsequent experiments, we use a fixed set of such parameters. Specifically, the dimensionality of the latent reference layer is set to $d=2$. The maximum number of iterations for sampling both degrees and community sizes is set to 1,0001{,}000. For the rewiring step, we use $0.05$ as the proportion of edges to be rewired in each batch, and $100$ as the number of rewiring batches.
It is worth noting that, like any artificial model, mABCD accepts only a feasible range of configuration values. Therefore, after retrieving the parameters for timik, we tailored them to fit within the constraints of the framework according to Section 4.2. The code to extract all parameters mentioned above from any network is also present in the previously mentioned GitHub repository, so even inexperienced user can generate a “digital twin” of any real network.
@@ 6.3 Experiments
Following this hypothesis, we generated four types of networks with modified values of $\xi$ as follows: series 2 with $\xi_{i}=1.00$, series 3 with $\xi_{i}=2\xi_{i_{\textit{timik}}}$, series 4 with $\xi_{i}=0.5\xi_{i_{\textit{timik}}}$, and series 5 with $\xi_{i}=0.01$. In other words, two types of networks exhibit extreme inter-community noise levels across all layers, while the remaining two have intermediate values, determined proportionally to those of the baseline. We then selected a spreading regime for the baseline networks (i.e., “digital twins” of timik, denoted as series 1). Specifically, we used the Neighbourhood Size Discount [czuba2024rankrefininginfmaxmln] to select seeds, and considered two sets of parameters for each protocol function: $(\dot{\delta},\dot{\pi},s)\in\{AND\}\times\{0.15,0.2,0.25,0.3\}\times\{1,3,5,7\}$ and $(\dot{\delta},\dot{\pi},s)\in\{OR\}\times\{0.03,0.05,0.07,0.09\}\times\{1,5,10,15\}$, where $\dot{\delta}$ and $\dot{\pi}$ denote the parameters of the MICM (see Section 6.1), while $s$ refers to the seed set budget, expressed as a percentage of the number of actors constituting a given network.
@@
With the second experiment, we aimed to address one of the fundamental questions in the domain of social network analysis: is it possible to compress a network in such a way that it preserves its topology while being smaller? Our goal was to examine this issue in the context of spreading phenomena. Specifically, we were interested in whether manipulating the number of actors, while preserving all other structural properties of the network (as dictated by the configuration model underlying mABCD), would result in different diffusion characteristics. It is worth noting that the approach adopted in this study can be considered novel, as most existing work reduces graphs using dedicated compression techniques rather than by explicitly employing standalone graph generators (for a comprehensive review, see [sun2024gcbench]).
@@
As one can note, for both protocol functions and across all evaluated activation probabilities, downsampling the network does not disturb spreading effectiveness (the differences are minimal). On the other hand, when the network gets upsampled, we observe slight, but non-negligible changes, in the diffusion — the more upsampled the network is, the higher $\Gamma$ change we observe. However, the disparities are not as big as in the results from Section 6.3. At this point, one might wonder why network downsampling does not alter the spreading trend, while upsampling does. A possible explanation lies in the average degree of the evaluated networks, as this parameter has a substantial impact on the effectiveness of diffusion under the MICM. We observe that downsampling has little effect on the average degree, whereas upsampling increases it by approximately $0.9$ (from around $20.2$ to $21.1$). Consequently, in the upsampled graphs, the diffusion process may be more effective due to this slight increase in connectivity, especially when the seed set is selected using a heuristic that prioritizes central actors rather than selects them at random.
@@
As in the previous experiments, the networks were also generated according to the configuration derived from timik. All of them had the $r$ value on the first layer set to $1.000$ in order to comply with the assumed seed selection scenario. However, they differed in the $r$ values on the remaining layers, as follows. Two series were generated to represent stronger community correlations: series 10 with $r_{2}=r_{3}=1.000$, and series 11 with $r_{2}=r_{3}=0.667$. Additionally, networks with weaker correlations were generated: series 12 with $r_{2}=r_{3}=0.333$, and series 13 with $r_{2}=r_{3}=0.001$. The spreading conditions applied were the same as in the previous experiments, with one difference in the seed selection method. Specifically, instead of employing a heuristic that considers actors’ representations across all network layers, we modified it to consider only the first layer (which is maximally correlated with the latent layer used in network generation). Consequently, the Neighbour Size Discount heuristic was reduced to its baseline version designed for single-layer networks: Degree Centrality Discount [Chen2009DegreeDiscount] and with that method a seed set was selected.
@@ 7 Conclusions
In this work, we have presented mABCD, a framework for generating multilayer networks with community structure and inter-layer relationships. We hope it addresses the existing gap in the field through its highly customizable configuration options and efficient algorithmic design. Furthermore, we release its implementation in Julia, along with Python ports, to make mABCD accessible to the broader research community. As mABCD builds upon the well-established generator introduced by [kaminski2021artificial], it can be regarded as a new member of the ABCD family. In addition to outlining the essential components of the model, we have validated its properties, demonstrating that mABCD produces stable results and outperforms its closest competitors in both efficiency and usage flexibility.
The paper also shows potential applications of the framework, illustrated in the context of studying spreading phenomena. Through three experiments, we examined whether and how selected parameters of the mABCD configuration model contribute to diffusion effectiveness. The results are nuanced rather than clear-cut. While there is a consistent trend with respect to the number of actors, the level of inter-community noise affects diffusion depending on the imposed regime. Finally, the strength of correlation between groups appears to have little influence on information spread. Altogether, these findings suggest that the proposed model is capable of supporting the analysis of spreading processes in multilayer networks, albeit up to certain extend.
As for the limitations of the proposed approach, the most significant one arises from the design assumptions underlying the model’s development. Specifically, mABCD is capable of generating scale-free networks, and therefore its applications are restricted to systems that conform to this property (e.g., social ones). Another constraint concerns the size of the generated network. Although it should not be regarded as the major shortcoming, it is important to note that mABCD fails to properly generate structures when certain parameters are assigned excessively low values, such as the minimum degree ($\delta_{i}$), the minimum community size ($s_{i}$), or the number of actors ($n$). The next limitation, which became particularly apparent during the experiments described in Section 6, concerns the difficulty of representing the properties of a real-world network within the parameter space of mABCD, especially the correlation strength between communities and the latent reference layer ($r_{i}$) which constitutes a core of the third phase in the network construction process (see Section 4.3). Although we have proposed a methodology for the coarse estimation of this value, a follow-up study is required to address this issue comprehensively.
Apart from the above, future work may proceed in two directions. The first concerns the ABCD model itself, which could be extended to other types of networks, such as temporal ones or multilayer ones without a lower bound imposed on the network size. The second way relates to the experiments involving spreading phenomena. In this context, it would be valuable to evaluate how other parameters of mABCD influence the dynamics of spreading processes. However, a more important and promising direction is the development of a metric for quantifying the difference between a real-world network and its “digital twin” generated by mABCD. Such a method could open a new avenue for influence maximization studies, as it would allow results and conclusions to be more directly related to real-world systems.
@@
This research was partially supported by the National Science Centre, Poland, under Grant no. 2022/45/B/ST6/04145999https://multispread.pwr.edu.pl/, the Polish Ministry of Education and Science within the programme “International Projects Co-Funded”, and the EU under the Horizon Europe, grant no. 101086321 OMINO. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the National Science Centre, Polish Ministry of Education and Science, the EU or the European Research Executive Agency.